OpenStudy (anonymous):

Find all solutions in the interval [0, 2pi). give exact values and show all algebra. 2cos^2 X +3sinX-3=0

OpenStudy (anonymous):

trick is to rewrite \[cos^2(x)= 1-sin^2(x)\]

OpenStudy (anonymous):

then get a quadratic equation in \[sin(x)\]

OpenStudy (anonymous):

\[2cos^2(x)+3sin(x)-3=0\] \[2(1-sin^2(x))+3sin(x)-3=0\] \[2-2sin^2(x)+3sin(x)-3=0\] \[-2sin^2(x)+3sin(x)-1=0\] \[2sin^2(x)-3sin(x)-1=0\]

OpenStudy (anonymous):

just solve as a quadratic. if it is easier to see replace sin(x) by z to get \[2z^2-3z+1=0\] \[(2z-1)(z-1)=0\] \[z=\frac{1}{2}\] or \[z=1\]

OpenStudy (anonymous):

then replace z by sin(x) and solve for x \[sin(x)=1\] \[x=\frac{\pi}{2}\]

OpenStudy (anonymous):

\[sin(x)=\frac{1}{2}\] \[x=\frac{\pi}{6}\] or \[x=\frac{5\pi}{6}\]

OpenStudy (anonymous):

ooops second answer is wrong sorry

OpenStudy (anonymous):

\[sin(x)=\frac{1}{2}\] \[x=\frac{11\pi}{6}\]

OpenStudy (anonymous):

k?

OpenStudy (anonymous):

so x=11pi/6?

OpenStudy (anonymous):

there are two places on the interval from 0 to 2pi where sine is 1/2

OpenStudy (anonymous):

one is at \[\frac{\pi}{6}\] and the other is at \[\frac{11\pi}{6}\] so all together you have 3 answers. those two and also \[\frac{\pi}{2}\] where sine is 1

OpenStudy (anonymous):

ohh i see thank you!

OpenStudy (anonymous):

welcome!

OpenStudy (anonymous):

any more?

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

any more of these to do?

OpenStudy (anonymous):

well i just put some up

OpenStudy (anonymous):

Does anyone happen to know 2cos^2 x+sinx-1=0 solve for x on [0,2pi)

OpenStudy (anonymous):

@satellite73